OpenStudy (anonymous):

what is the vaule of the expression -(3x -2)+x+9 when x=-3...?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

-(9+2)+3+9 -(7)+3+9 -7+12 =5

OpenStudy (anonymous):

wait you said -3 hold on

OpenStudy (anonymous):

-(-11)+-3+9 11-3+9 8+9 =17

OpenStudy (anonymous):

thanks butbut it dosent have 17 on my papper

OpenStudy (anonymous):

what thats preposterous!!! hahah hmmmm

OpenStudy (anonymous):

are you sure you wrote down the problem right?

OpenStudy (anonymous):

lol yeaa whatt is the vaule of the expression -2(3x-2)+x+9 when x=-3

OpenStudy (anonymous):

see you missed the first 2 when you wrote it the first tim ehaha hold on

OpenStudy (anonymous):

-2(-9-2)+-3+9 -2(-11)-3+9 22-3+9 19+9 =27 is that on your sheet ?